Four point-like charges are placed as shown in the
Answer #1

here

q1 = 15 uC

q1 = 15 * 10^-6 C

q2 = -28 uC

q2 = -28 * 10^-6 C

q3 = + 24 uC

q3 = 24 * 10^-6 C

q4 = - 32 uC

q4 = -32 * 10^-6 C

a = 26 cm

a = 0.26 m

the electric potential energy of the entire system is ,Utotal = U12 + u13 + U14 + U23 + U24 + U34

Utotal = 9*10^9 * 10^-6 * 10^-6 *(15 * (-28)/.26 - 15 * 32/(sqrt(2) * 0.13) + 24 * 15 /(sqrt(2) * 0.26) - 24 * 28/0.26 + 32 * 28/(sqrt(2) * 0.13) - 32 * 24/(sqrt(2) * 0.13) )

Utotal = - 46.22 J

the electric potential energy of the entire system is - 46.22 J